OpenStudy (anonymous):

subtract 2/49z^3y - 1/14z^2y

OpenStudy (anonymous):

find the common factor?

OpenStudy (kropot72):

Use as a common denominator: \[(49\times14)z ^{3}y\] Then combining the two terms of the question: \[(14-49z)\div((49\times14)z ^{3}y)=(2-7z)\div(98z ^{3}y)\]

jhonyy9 (jhonyy9):

2 1 2*2 - 1*7 4 - 7 -3 ------- - -------- = ----------- = -------- = -------- 49z^3y 14z^2y 98z^3y 98z^3y 98z^3y